An electron with an initial speed of 660,000 m/s is brought to rest by an electric field.

What was the potential difference that stopped the electron?

What was the initial kinetic energy of the electron, in electron volts?
Answer #1

Change in kinetic energy = k2 – k1

= 0 – k1

= - 0.5mv2

= - 0.5(9.1×10-31)(5×105)2

∆k = - 1.375 × 10-19 joule

& 1ev = 1.6 ×10-19joule

Change in KE = - 0.86 ev

& change in potential energy= - change in kinetic energy

= 1.375× 10-19 joule

Change in potential=change in potential energy/q

=( 1.375 × 10-19)/( 1.6 ×10-19)

= 0.86 volt

K1=0.86v